OpenStudy (anonymous):

Which ratio is not equal to 4/9? A.45/95 B.16/36 C.8/18 D.20/45

OpenStudy (cwrw238):

well the numerator must be a multiple of 4 and denominator a multiple of 9.

OpenStudy (anonymous):

C. 8/18?

OpenStudy (cwrw238):

no because 8 is 2*4 and 18 = 2*9

OpenStudy (cwrw238):

what you are looking for is a fraction which does not fit these requirements

OpenStudy (cwrw238):

sorry - i might have confused you

OpenStudy (anonymous):

Just a bit... >.<~

OpenStudy (cwrw238):

20/45 = 4/9 ( = 5*4 / 5*9) so its not that one

OpenStudy (anonymous):

Personally I THINK it's the first one. But that's just an educated guess because those number don't seem very common.

OpenStudy (cwrw238):

yes - right - simply because 45 is not a multiple of 4 - neither is 95 a multiple of 9